Two men can paint a fence in X hours and (X + 1) hours respectively. Find the time it will take the two men to paint the fence together.
Answer by stanbon(75887)

Two men can paint a fence in X hours and (X + 1) hours respectively. Find the time it will take the two men to paint the fence together.
------
1st man rate:: 1/x job/hr
2nd man rate:: 1/(x+1) job/hr
----
Together rate: 1/k job/hr
------
Equation:
rate + rate = together rate
1/x + 1/(x+1) = 1/k
-----
k(x+1) + k(x) = x(x+1)
-----
k[x+1+x] = x^2+x
----
k = (x^2+x)/(2x+1) hrs (time to do the job together)
==================
Cheers,
Stan H.